# Public-Key Quantum Money and Fast Real Transforms

### v3

@@ 1 Introduction
In a seminal paper, Wiesner [29] introduced the concept of quantum money, where bills are represented by quantum states. In constracst with classical money, quantum money cannot be couterfeited using general copying machines. This is a consequence of a fundamental theorem in quantum mechanichs called the no-cloning theorem. Wiesner’s scheme, which in is in modern lterminology called a private-key quantum money scheme, had siginificant drawbacks. In particular, in a private-key scheme the bank is required to verify each bill, meaning it must be involved in every transaction. This is the mean reason that these schemes are not generally practical.
In 2009, Aaronson [1] propoased the first concrete proposal for a different type of quantum money, known as public-key quantum money, that did not have such drawbacks. In publick-key quantum money anyone can verify the bill while only the bank can issue it. Aaronson’s scheme was later broken by Lutomirski et al. [24]. In the years since, several alternative constructions have been explored [2, 17, 30, 19, 20, 23, 31], yet each has either been broken [14, 26, 7, 23] or relies on non-standard cryptographic assumptions.
@@ 2 Preliminaries
[…]
@@ 2.1 Group actions
For a group $G$ and a set $X$, we say that $G$ acts on $X$ if there is a mapping $*:G\times X\to X$ that satisfies the following properties:
@@
    Compatibility: for every $a,b\in G$ and every $x\in X$, $g*(h*x)=(gh)*x$,
@@
    Identity: for the identity $1\in G$ and every $x\in X$, $1*x=x$.
We use the notation $(G,X,*)$ to denote a group $G$ acting on a set $X$ through the action *. A group action is called regular if for every $x,y\in X$ there exists a unique $g\in G$ such that $g*x=y$. In order for a group action $(G,X,*)$ to be suitable for algorithmic applications, the group $G$, the set $X$, and the action * must adhere to specific properties. This motivates the concept of an effective group action.
Let $G$ be a finite group and $X$ a finite set. A group action $(G,X,*)$ is said to be effective if it satisfies the following properties: i) There are efficient algorithms for elementary operations such as membership testing, equality testing, sampling uniform elements and group operation for $G$, ii) There are efficient algorithms for membership testing and unique representation in $X$, and iii) There is exists an efficient algorithm for the action *.
@@
Let $(G,X,*)$ be an effective group action. Given a pair $(x,g*x)$, where $g\in G$, the discrete logarithm problem (DLP) is to compute $g$.
We say that the DLP assumption holds for $(G,X,*)$ if no quantum polynomial time (QPT) algorithm can solve the DLP with respect to $(G,X,*)$. A group action for which the DLP assumption holds is referred to as a cryptographic group action.
@@ 2.2 The Fourier transform
The quantum Fourier transform of a (normalized) state $\sum_{g\in G}f(g)\lvert g\rangle$ is given by $\sum_{x\in G}\hat{f}(x)\lvert x\rangle$. For a regular group action $(G,X,*)$, any subset $S\subseteq G$, any $y\in X$, and any $h\in G$, we define
@@
These state are simultaneous eigenstates of the group action operation. Specifically, for the unitary $U_{k}:\lvert y\rangle\mapsto\lvert k*y\rangle$, where $k\in G$, we have $U_{k}\lvert G^{(h)}*x\rangle=\chi(-k,h)\lvert G^{(h)}*x\rangle$. These states resemble the set of Fourier states over the abelian group $G$. We will also sometimes refer to them as Fourier states.
@@ 3 Fast Real Transforms
where we have skipped the normalization for the left hand side state. Here, $\mathsf{H}_{G}(f)$ can be either of the transforms (2) or (4). For a single basis element of the cyclic group $\mathbb{Z}_{N}$, the quantum Hartly transform simplifies to
@@
The quantum version of this transform is define by
@@ 4 A New Algorithm for $\mathsf{QHT}$
We now show how to compute $\mathsf{QHT}_{N}$ recursively. Given $a\in\mathbb{Z}_{N}$, we once again express $\lvert a\rangle=\lvert t\rangle\lvert b\rangle$, where $b$ is the least significant bit, so that $a=2t+b$ for some $t\in\mathbb{Z}_{N/2}$. Suppose we have already have an efficient circuit for computing $\mathsf{QHT}_{N/2}$. To compute the Hartley transform of $\lvert a\rangle$, we introduce an ancilla qubit, resulting in the state $\lvert 0\rangle\lvert t\rangle\lvert b\rangle$. Then
@@
[…]
@@
    Initialize an ancilla qubit to 00 to obtain the state $\lvert 0\rangle\lvert\psi\rangle$
@@
Algorithm 1 is correct and can be implemented using $\approx 2\log^{2}N+O(\log N)$ elementary gates.
@@
for $j=0,1,\dots,n-1$. For $b=0$, $R(y,0)=\mathds{1}$, and for $b=1$, $R(y,1)$ is the product of the $R_{j}$, where the control qubit is the $j$the qubit of $y$. Therefore, for a $y$ of size $k=\lceil\log y\rceil$, we can construct $U_{R}$ using $k$ gates form the set $\{R_{j}\}$.
Let $T(N)$ be the gate complexity of Algorithm 1 for an input state of dimension $N$. Assuming access to the gates $R_{j}$ as elementary gates, it follows from above that $T(N)\approx T(N/2)+4\log N+O(1)$. Therefore, $T(N)\approx 2\log^{2}N+O(\log N)$. ∎
@@ 4.1 Comparison With Other Algorithms
    Compute a controlled two’s complement of a $(n-2)$-bit integer
@@
[…]
@@
    Compute a controlled two’s complement of a $(n-2)$-bit integer
Each of the above steps require approximately $\lceil\log N\rceil$ elementary gates to implement. Let $T(N)$ denote the gate complexity of the algorithm for input size $N$. Then $T(N)\approx T(N/2)+5\log N+O(1)$, which implies that
$T(N)\approx\frac{5}{2}\log^{2}N+O(\log N).$
@@
$T(N)\approx\frac{5}{2}\log^{2}N+O(\log N).$
It follows from the above and Theorem 4.1 that our algorithm for $\mathsf{QHT}_{N}$ has a gate complexity that is approximately 1.25×1.25$\times$1.25 × lower than that of [3] and [22].
@@ 5 Other Real Transforms
where $k_{j}=1/\sqrt{2}$ for $j=N$ and $k_{j}=1$ for $j\neq N$. Different version of the discrete cosine transform are
@@
The quantum sine transform $\mathsf{QS}_{N-1}^{\text{I}}$ can be implemented using $\approx\log^{2}N+O(\log N)$ elementary gates.
@@ 6.1 Quantum Money From Group Actions
  • $\mathsf{Gen}(1^{\lambda})$: This algorithm takes a security parameter $\lambda$ as input and outputs a pair $(s,\rho_{s})$, where $s$ is a binary string called the serial number, and $\rho_{s}$ is a quantum state called the banknote. The pair $(s,\rho_{s})$, or simply $\rho_{s}$, is sometimes denoted by \$currency-dollar\\$\$.
  • $\mathsf{Ver}(s,\rho_{s})$: This algorithm takes a serial number and an alleged banknote as input and outputs either $1$ (accept) or 00 (reject).
@@
  • $\mathsf{Ver}(h,\lvert\psi\rangle)$. First, check whether $\lvert\psi\rangle$ has support in $X_{\lambda}$. If not, return 00. Then, apply $\mathsf{cmpIndex}$ to the state $\lvert\psi\rangle\lvert 0\rangle$, and measure the second register to obtain some $h^{\prime}\in G_{\lambda}$. If $h^{\prime}=h$, return $1$; otherwise return 00.
@@ 7 Quantum Money With The Hartley Transform
  • $\mathsf{Ver}(h,\lvert\psi\rangle)$. First, check whether $\lvert\psi\rangle$ has support in $X$. If not, return 00. Then, apply $\mathsf{cmpIndex}$ to the state $\lvert 0\rangle\lvert\psi\rangle$, and measure the first register to obtain some $h^{\prime}\in\mathbb{Z}_{N}$. If $h^{\prime}=h$, return $1$; otherwise return 00.
@@
Unfortunately, we do not how to use the resulting state in (14) to verify the banknote $(h,\lvert\mathbb{Z}_{N}^{(h)}*x\rangle_{H})$. In particular, when given the banknote $(h,\lvert\mathbb{Z}_{N}^{(-h)}*x\rangle_{H})$, the above verification algorithm produces the state
@@ 7.1 Group Actions with Twists
We show how twists can be used to successfully verify banknotes in the above money scheme. A group action $(G,X,*)$, with a fixed element $x\in X$, is said to support twists if there exists an efficient algorithm that computes the mapping $g*x\mapsto(-g)*x$ for any $g\in G$. The twisting map can be efficiently performed in isogeny-based group actions, which are the most widely used cryptographic group actions. We assume that the group actions used to construct the above quantum money scheme supports twists.
@@
Therefore, applying $T_{u}$ results in one of these states, depending on the input state. Both of these states are eigenstates of the twist unitary. The former corresponds to eigenvalue $1$ and latter corresponds to eigenvalue $-1$.
@@
    If the support of $\lvert\psi\rangle$ is not in $X$, return 00.
@@
    If the measurement outcome is $1$, return 00.
@@
    Use Lemma 7.2 to check whether the resulting state is $\lvert\mathbb{Z}_{N}^{(h)}*x\rangle$. If yes, return $1$; otherwise, return 00.
Some explanations about the algorithm are in order. Step 1 can be implemented using an extra qubit initialized to $\lvert 0\rangle$, which is set to $1$ if the values of the first register is no in $X$. This qubit is then measured, and the state is rejected if the measurement outcome is not 00. The post-measurement state, which we denote again by $\lvert\psi\rangle$, has support in $X$. Similarly, implementing the measurement in Step 3 can be done efficiently, as $h$ is known. Finally, the last step is implemented using an extra qubit, as stated Lemma 7.2.
@@
The final step of the algorithm projects this state onto $\lvert\mathbb{Z}_{N}^{(h)}*x\rangle_{H}$ with probability $\lvert\alpha_{h}^{2}\rvert/(\lvert\alpha_{h}\rvert^{2}+\lvert\alpha_{-h}^{2}\rvert)$. Therefore, the probability of the state $\lvert\psi\rangle$ passing verification is
@@ 8 Computing The Serial Number Using Quantum Walks
In this section, we show how to use continuous-time quantum walks to compute the serial number of a given money state. Quantum walks are quantum analogs of classical random walks and play a fundamental role in quantum algorithms. Similar to the classical case, there are two types of quantum walks: continuous-time and discrete-time, both of which exhibit significantly different behaviours compared to classical random walks. In particular, a key distinction between quantum and classical random walks lies in the wave nature of quantum mechanics. In a quantum walk, interference effects can occur, allowing computational advantage for exploration of the graph compared to classical walks, e.g., [4, 5, 10, 11].
@@
A discrete-time quantum walk on $\Gamma$ can be described using the adjacency matrix $A$ by making $A$ into a stochastic matrix [27]. More precisely, we define a matrix $P$ as $P_{jk}=A_{jk}/\deg(j)$, where $\deg(j)$ is the degree of the vertex $v_{j}$ in $\Gamma$. If the $\Gamma$ has $N$ vertices, the discrete time quantum walk on $\Gamma$ is defined by a unitary operator on the finite Hilbert space $\mathbb{C}^{N}\times\mathbb{C}^{N}$ as follows. Define the states
@@
[…]
@@
The discrete-time quantum walk is then defined by $W=iS(2TT^{*}-\mathds{1})$, where $S$ is the swap operator acting as $S\lvert j_{1}\rangle\lvert b_{1}\rangle\lvert j_{2}\rangle\lvert b_{2}\rangle=\lvert j_{2}\rangle\lvert b_{2}\rangle\lvert j_{1}\rangle\lvert b_{1}\rangle$ for all $10\leq j_{1},j_{2}\leq N-$ 1 and $b_{1},b_{2}\in\{0,1\}$. To efficiently simulate the continuous-time quantum walk defined by the Hamiltonian $H$, we need to be able to efficiently apply the isometry $T$ and its inverse $T^{*}$, as well as the walk operator $W$. Assuming black-box query access to $H$, we have the following:
@@
Let $H$ be a $d$-sparse Hamiltonian $H$ acting on $n$ qubits. Then, the unitary $e^{-iHt}$ can be approximated to withing error $\epsilon$ using
@@
8.1 Group Action Quntum Walks
Let $G$ be an abelian group and let $Q=\{q_{1},q_{2},\dots,q_{k}\}\subset G$ be a symmetric set, i.e., $q\in Q$ if and only if $-q\in Q$. The Cayley graph associated to $G$ and $Q$ is a graph $\Gamma=(V,E)$, where the vertex set is $V=G$, and the edge set $E$ consists of pairs $(a,b)\in G\times G$ such that there exists $q\in Q$ with $b=q+a$. The adjacency matrix of $\Gamma$ can be expressed as

### v4

@@ 1 Introduction
In a seminal paper, Wiesner [29] introduced the concept of quantum money, where bills are represented by quantum states. In contrast with classical money, quantum money cannot be counterfeited using general copying machines. This is a consequence of a fundamental theorem in quantum mechanics called the no-cloning theorem. Wiesner’s scheme, which in modern terminology is called a private-key quantum money scheme, had significant drawbacks. In particular, in a private-key scheme the bank is required to verify each bill, meaning it must be involved in every transaction. This is the main reason that these schemes are not generally practical.
In 2009, Aaronson [1] proposed the first concrete proposal for a different type of quantum money, known as public-key quantum money, that did not have such drawbacks. In public-key quantum money anyone can verify the bill while only the bank can issue it. Aaronson’s scheme was later broken by Lutomirski et al. [24]. In the years since, several alternative constructions have been explored [2, 17, 30, 19, 20, 23, 31], yet each has either been broken [14, 26, 7, 23] or relies on non-standard cryptographic assumptions.
@@ 2 Preliminaries
We follow the presentation of [15] regarding group actions and quantum computation throughout this paper. A finite Hilbert space $\mathcal{X}$ of dimension $N$ is a complex Euclidean space, isomorphic as a $\mathbb{C}$-vector space to $\mathbb{C}^{N}$. An $N$-dimensional quantum state is represented by a unit vector $\lvert\psi\rangle\in\mathcal{H}$, where $\lvert\cdot\rangle$ denotes Dirac notation. In practice, we typically use specific bases for a Hilbert space. For example, when considering a finite group $G$, we work with the Hilbert space $\mathcal{X}=\mathbb{C}^{G}$, which is a $\lvert G\rvert-$dimensional space spanned by the basis $\{\lvert g\rangle:g\in G\}$.
@@ 2.1 Group actions
For a group $G$ and a set $X$, we say that $G$ acts on $X$ if there is a mapping $G\times X\to X$ that satisfies the following properties:
@@
    Compatibility: for every $g,h\in G$ and every $x\in X$, g(hx)=(gh)xg*(h*x)=(gh)*x,
@@
    Identity: for the identity $1\in G$ and every $x\in X$, 1x=x1*x=x.
We use the notation (G,X,)(G,X,*) to denote a group $G$ acting on a set $X$ through the action *. A group action is called regular if for every $x,y\in X$ there exists a unique $g\in G$ such that gx=yg*x=y. In order for a group action (G,X,)(G,X,*) to be suitable for algorithmic applications, the group $G$, the set $X$, and the action * must adhere to specific properties. This motivates the concept of an effective group action.
Let $G$ be a finite group and $X$ a finite set. A group action (G,X,)(G,X,*) is said to be effective if it satisfies the following properties: i) There are efficient algorithms for elementary operations such as membership testing, equality testing, sampling uniform elements and group operation for $G$, ii) There are efficient algorithms for membership testing and unique representation in $X$, and iii) There exists an efficient algorithm for the action *.
@@
Let (G,X,)(G,X,*) be an effective group action. Given a pair (x,gx)(x,g*x), where $g\in G$, the discrete logarithm problem (DLP) is to compute $g$.
We say that the DLP assumption holds for (G,X,)(G,X,*) if no quantum polynomial time (QPT) algorithm can solve the DLP with respect to (G,X,)(G,X,*). A group action for which the DLP assumption holds is referred to as a cryptographic group action.
@@ 2.2 The Fourier transform
The quantum Fourier transform of a (normalized) state $\sum_{g\in G}f(g)\lvert g\rangle$ is given by $\sum_{x\in G}\hat{f}(x)\lvert x\rangle$. For a regular group action (G,X,)(G,X,*), any subset $S\subseteq G$, any $y\in X$, and any $h\in G$, we define
@@
These states are simultaneous eigenstates of the group action operation. Specifically, for the unitary $U_{k}:\lvert y\rangle\mapsto\lvert k*y\rangle$, where $k\in G$, we have $U_{k}\lvert G^{(h)}*x\rangle=\chi(-k,h)\lvert G^{(h)}*x\rangle$. These states resemble the set of Fourier states over the abelian group $G$. We will also sometimes refer to them as Fourier states.
@@ 3 Fast Real Transforms
where we have skipped the normalization for the left hand side state. Here, $\mathsf{H}_{G}(f)$ can be either of the transforms (2) or (4). For a single basis element of the cyclic group $\mathbb{Z}_{N}$, the quantum Hartley transform simplifies to
@@
The quantum version of this transform is defined by
@@ 4 A New Algorithm for $\mathsf{QHT}$
We now show how to compute $\mathsf{QHT}_{N}$ recursively. Given $a\in\mathbb{Z}_{N}$, we once again express $\lvert a\rangle=\lvert t\rangle\lvert b\rangle$, where $b$ is the least significant bit, so that $a=2t+b$ for some $t\in\mathbb{Z}_{N/2}$. Suppose we already have an efficient circuit for computing $\mathsf{QHT}_{N/2}$. To compute the Hartley transform of $\lvert a\rangle$, we introduce an ancilla qubit, resulting in the state $\lvert 0\rangle\lvert t\rangle\lvert b\rangle$. Then
@@
$\lvert\psi\rangle$ $\mapsto\frac{1}{\sqrt{N/2}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big(\frac{2\pi ay}{N}\Big)\lvert 0\rangle\lvert y\rangle\lvert b\rangle$ $(\textsc{cnot})$
@@
    Initialize an ancilla qubit to 0 to obtain the state $\lvert 0\rangle\lvert\psi\rangle$
@@
Algorithm 1 is correct and can be implemented using $2\log^{2}N+O(\log N)$ elementary gates.
@@
for $j=0,1,\dots,n-1$. For $b=0$, $R(y,0)=\mathds{1}$, and for $b=1$, $R(y,1)$ is the product of the $R_{j}$, where the control qubit is the $j$th qubit of $y$. Therefore, for a $y$ of size $k=\lceil\log y\rceil$, we can construct $U_{R}$ using $k$ gates from the set $\{R_{j}\}$.
Let $T(N)$ be the gate complexity of Algorithm 1 for an input state of dimension $N$. Assuming access to the gates $R_{j}$ as elementary gates, it follows from above that $T(N)=T(N/2)+4\log N+O(1)$. Therefore, $T(N)=2\log^{2}N+O(\log N)$. ∎
@@ 4.1 Comparison With Other Algorithms
    Compute a controlled two’s complement of a (n2)(n-2)-bit integer
@@
    Compute the multi-controlled phase gate $\lvert 1x\rangle\mapsto-\lvert 1x\rangle$, $\lvert 10\rangle\mapsto\lvert 10\rangle$.
@@
    Compute a controlled two’s complement of a (n2)(n-2)-bit integer
Each of the above steps requires approximately $\lceil\log N\rceil$ elementary gates to implement. Let $T(N)$ denote the gate complexity of the algorithm for input size $N$. Then $T(N)=T(N/2)+5\log N+O(1)$, which implies that
$T(N)=\frac{5}{2}\log^{2}N+O(\log N).$
@@
$T(N)=\frac{5}{2}\log^{2}N+O(\log N).$
It follows from the above and Theorem 4.1 that our algorithm for $\mathsf{QHT}_{N}$ has a gate complexity that is approximately $1.25\times$ lower than that of [3] and [22].
@@ 5 Other Real Transforms
where $k_{j}=1/\sqrt{2}$ for $j=N$ and $k_{j}=1$ for $j\neq N$. Different versions of the discrete cosine transform are
@@
The quantum sine transform $\mathsf{QS}_{N-1}^{\text{I}}$ can be implemented using $2\log^{2}N+O(\log N)$ elementary gates.
@@ 6.1 Quantum Money From Group Actions
[…]
  • $\mathsf{Ver}(s,\rho_{s})$: This algorithm takes a serial number and an alleged banknote as input and outputs either $1$ (accept) or 0 (reject).
@@
  • $\mathsf{Ver}(h,\lvert\psi\rangle)$. First, check whether $\lvert\psi\rangle$ has support in $X_{\lambda}$. If not, return 0. Then, apply $\mathsf{cmpIndex}$ to the state $\lvert\psi\rangle\lvert 0\rangle$, and measure the second register to obtain some $h^{\prime}\in G_{\lambda}$. If $h^{\prime}=h$, return $1$; otherwise return 0.
@@ 7 Quantum Money With The Hartley Transform
  • $\mathsf{Ver}(h,\lvert\psi\rangle)$. First, check whether $\lvert\psi\rangle$ has support in $X$. If not, return 0. Then, apply $\mathsf{cmpIndex}$ to the state $\lvert 0\rangle\lvert\psi\rangle$, and measure the first register to obtain some $h^{\prime}\in\mathbb{Z}_{N}$. If $h^{\prime}=h$, return $1$; otherwise return 0.
@@
Unfortunately, we do not know how to use the resulting state in (14) to verify the banknote $(h,\lvert\mathbb{Z}_{N}^{(h)}*x\rangle_{H})$. In particular, when given the banknote $(h,\lvert\mathbb{Z}_{N}^{(-h)}*x\rangle_{H})$, the above verification algorithm produces the state
@@ 7.1 Group Actions with Twists
We show how twists can be used to successfully verify banknotes in the above money scheme. A group action (G,X,)(G,X,*), with a fixed element $x\in X$, is said to support twists if there exists an efficient algorithm that computes the mapping $g*x\mapsto(-g)*x$ for any $g\in G$. The twisting map can be efficiently performed in isogeny-based group actions, which are the most widely used cryptographic group actions. We assume that the group actions used to construct the above quantum money scheme support twists.
@@
Therefore, applying $T_{u}$ results in one of these states, depending on the input state. Both of these states are eigenstates of the twist unitary. The former corresponds to eigenvalue $1$ and the latter corresponds to eigenvalue 1-1.
@@
    If the support of $\lvert\psi\rangle$ is not in $X$, return 0.
@@
    If the measurement outcome is $1$, return 0.
@@
    Use Lemma 7.2 to check whether the resulting state is $\lvert\mathbb{Z}_{N}^{(h)}*x\rangle$. If yes, return $1$; otherwise, return 0.
Some explanations about the algorithm are in order. Step 1 can be implemented using an extra qubit initialized to $\lvert 0\rangle$, which is set to $1$ if the values of the first register is not in $X$. This qubit is then measured, and the state is rejected if the measurement outcome is not 0. The post-measurement state, which we denote again by $\lvert\psi\rangle$, has support in $X$. Similarly, implementing the measurement in Step 3 can be done efficiently, as $h$ is known. Finally, the last step is implemented using an extra qubit, as stated in Lemma 7.2.
@@
[…]
@@ 8 Computing The Serial Number Using Quantum Walks
In this section, we show how to use continuous-time quantum walks to compute the serial number of a given money state. Quantum walks are quantum analogs of classical random walks and play a fundamental role in quantum algorithms. Similar to the classical case, there are two types of quantum walks: continuous-time and discrete-time, both of which exhibit significantly different behaviors compared to classical random walks. In particular, a key distinction between quantum and classical random walks lies in the wave nature of quantum mechanics. In a quantum walk, interference effects can occur, allowing computational advantage for exploration of the graph compared to classical walks, e.g., [4, 5, 10, 11].
@@
A discrete-time quantum walk on $\Gamma$ can be described using the adjacency matrix $A$ by making $A$ into a stochastic matrix [27]. More precisely, we define a matrix $P$ as $P_{jk}=A_{jk}/\deg(j)$, where $\deg(j)$ is the degree of the vertex $v_{j}$ in $\Gamma$. If the $\Gamma$ has $N$ vertices, the discrete-time quantum walk on $\Gamma$ is defined by a unitary operator on the finite Hilbert space $\mathbb{C}^{N}\times\mathbb{C}^{N}$ as follows. Define the states
@@
$\lvert\phi_{j1}\rangle$ $:=\lvert 0\rangle\lvert 1\rangle,$
@@
The discrete-time quantum walk is then defined by $W=iS(2TT^{*}-\mathds{1})$, where $S$ is the swap operator acting as $S\lvert j_{1}\rangle\lvert b_{1}\rangle\lvert j_{2}\rangle\lvert b_{2}\rangle=\lvert j_{2}\rangle\lvert b_{2}\rangle\lvert j_{1}\rangle\lvert b_{1}\rangle$ for all $0\leq j_{1},j_{2}\leq N-1$ and $b_{1},b_{2}\in\{0,1\}$. To efficiently simulate the continuous-time quantum walk defined by the Hamiltonian $H$, we need to be able to efficiently apply the isometry $T$ and its inverse $T^{*}$, as well as the walk operator $W$. Assuming black-box query access to $H$, we have the following:
@@
Let $H$ be a $d$-sparse Hamiltonian $H$ acting on $n$ qubits. Then, the unitary $e^{-iHt}$ can be approximated to within error $\epsilon$ using
@@
8.1 Group Action Quantum Walks
Let $G$ be an abelian group and let $Q=\{q_{1},q_{2},\dots,q_{k}\}\subset G$ be a symmetric set, i.e., $q\in Q$ if and only if $q\in Q$. The Cayley graph associated to $G$ and $Q$ is a graph $\Gamma=(V,E)$, where the vertex set is $V=G$, and the edge set $E$ consists of pairs $(a,b)\in G\times G$ such that there exists $q\in Q$ with $b=q+a$. The adjacency matrix of $\Gamma$ can be expressed as